OpenStudy (anonymous):

Solve for x. (3x/4)+(2x/3)=51

OpenStudy (anonymous):

9x + 8x ---------- = 51 12 17x = 51 * 12 51 * 12 x = -------- = 36 17

OpenStudy (aroub):

ok u multiply by the lcd which is 12 so it becomes 9x+8x=612 , 17x=612 x=36